OpenStudy (anonymous):

7!.8!.3/5!.9!= solutions 282,240 21 14

OpenStudy (jdoe0001):

\(\bf 9!\implies 9\cdot 8\cdot 7\cdot 6\cdot 5\cdot 4\cdot 3\cdot 2\cdot 1\\ 8!\implies 8\cdot 7\cdot 6\cdot 5\cdot 4\cdot 3\cdot 2\cdot 1\\ 7!\implies 7\cdot 6\cdot 5\cdot 4\cdot 3\cdot 2\cdot 1\\ 5!\implies 5\cdot 4\cdot 3\cdot 2\cdot 1\) what do you think?

OpenStudy (anonymous):

i had 21